rssLink RSS for all categories
 
icon_red
icon_green
icon_red
icon_red
icon_blue
icon_green
icon_green
icon_red
icon_red
icon_green
icon_orange
icon_green
icon_green
icon_green
icon_red
icon_blue
icon_red
icon_orange
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_red
icon_orange
icon_green
 

FS#93 — FS#4201 — th2-1-6k

Attached to Project— Network
Incident
the whole network
CLOSED
100%
We will remove a card 6148 (ouff), put a new one 6407
and switch the 10G ports on this card in order to remove the card
7 which caused some problems recently.
Date:  Friday, 21 May 2010, 12:11PM
Reason for closing:  Done
Comment by OVH - Friday, 21 May 2010, 11:49AM

We remove the card 5.


Comment by OVH - Friday, 21 May 2010, 11:50AM

.May 21 09:42:48 GMT: %C6KERRDETECT-SP-2-SWBUSSTALL: The switching bus is experiencing stall for 3 seconds
.May 21 09:43:13 GMT: %C6KERRDETECT-SP-2-SWBUSSTALL_RECOVERED: The switching bus stall is recovered and data traffic switching continues

A small freeze of the bus.


Comment by OVH - Friday, 21 May 2010, 11:55AM

The links were switched on the new card.

We will remove the card 7.


Comment by OVH - Friday, 21 May 2010, 11:56AM

We cut off the ports on the card 5.


Comment by OVH - Friday, 21 May 2010, 12:01PM

May 21 11:13:20 80G.th2-1-6k.routers.ovh.net 162960: .May 21 10:13:04 GMT: %IPC-DFC5-5-WATERMARK: 879 messages pending in rcv for the port Card5/0:Request(2050000.4) seat 2050000
May 21 11:13:53 80G.th2-1-6k.routers.ovh.net 162961: .May 21 10:13:37 GMT: %IPC-DFC5-5-WATERMARK: 913 messages pending in rcv for the port Card5/0:Request(2050000.4) seat 2050000
May 21 11:14:26 80G.th2-1-6k.routers.ovh.net 162962: .May 21 10:14:11 GMT: %IPC-DFC5-5-WATERMARK: 950 messages pending in rcv for the port Card5/0:Request(2050000.4) seat 2050000

A human mistake during the card mounting: the RAM
strip has not been upgraded.

It is done.


Comment by OVH - Friday, 21 May 2010, 12:02PM

The ports were put back in use again.

Everything is ok.


Comment by OVH - Friday, 21 May 2010, 12:11PM

The sessions with Neuf are being restarted.


Comment by OVH - Friday, 21 May 2010, 12:11PM

The intervention is finalized.